# Reconstruction of boosted and resolved multi-Higgs-boson events with symmetry-preserving attention networks

### v1

@@ 6 Summary
We adapted the enhanced symmetry-preserving attention network (SPA-Net) to reconstruct multiple Higgs bosons in $\mathrm{HHH}$ and $\mathrm{HH}$ events, addressing both resolved and boosted topologies. Our $\mathrm{HHH}$ boosted+resolved SPA-Net training demonstrated superior Higgs boson reconstruction purity and efficiency compared to a baseline approach combining a $\chi^{2}$-based methods and a boosted decision tree (BDT) large-radius jet tagger, as shown in Table 3. In particular, for $\mathrm{HHH}$ ($\mathrm{HH}$) events considering both boosted and resolved topologies, our SPA-Net approach increases the $\mathrm{H}$ purity by 57% (62%) and the $\mathrm{H}$ efficiency by 38% (23%) compared to the baseline method.
@@
$\mathrm{H}$ Purity 0.251 0.430 0.260 0.408 0.390 0.702 0.402 0.651
$\mathrm{H}$ Efficiency 0.394 0.448 0.367 0.506 0.578 0.769 0.547 0.674
Reco. Target $\mathrm{H}$ 32,223 36,630 30,787 42,457 47,039 62,562 45,481 56,056
@@
In addition, we proposed using predictions of SPA-Net to construct a likelihood for the number of reconstructed boosted and resolved Higgs bosons. By further prioritizing boosted Higgs boson candidates and removing the overlapping resolved Higgs boson candidates, we can uniquely categorize events into boosted or resolved topologies.
Overall, our results show that the boosted+resolved SPA-Net algorithm provides significant improvements in Higgs boson reconstruction across both $\mathrm{HHH}$ and $\mathrm{HH}$ events, demonstrating its potential as a powerful tool for analyzing multi-Higgs boson final states.

### v3

@@ 6 Summary
We generalized the symmetry-preserving attention network (SPA-Net) to reconstruct multiple Higgs bosons in $\mathrm{HHH}$ and $\mathrm{HH}$ events in both resolved and boosted topologies simultaneously, improving reconstruction performance in both. Our $\mathrm{HHH}$ boosted+resolved SPA-Net training demonstrated superior Higgs boson reconstruction purity and efficiency compared to a baseline approach combining a $\chi^{2}$-based methods and a boosted decision tree (BDT) large-radius jet tagger, as shown in Table 3. In particular, for $\mathrm{HHH}$ ($\mathrm{HH}$) events considering both boosted and resolved topologies, our SPA-Net approach increases the $\mathrm{H}$ purity by 56% (80%) and the $\mathrm{H}$ efficiency by 38% (37%) compared to the baseline method.
@@
$\mathrm{H}$ Purity 0.251 0.444 0.261 0.409 0.390 0.702 0.402 0.719
$\mathrm{H}$ Efficiency 0.394 0.464 0.367 0.507 0.578 0.769 0.547 0.752
Reco. Target $\mathrm{H}$ 32,223 37,922 30,787 42,486 47,039 62,562 45,481 62,522
@@
In addition, we proposed using predictions of SPA-Net to construct a likelihood for the number of reconstructed boosted and resolved Higgs bosons. By further prioritizing boosted Higgs boson candidates and removing the overlapping resolved Higgs boson candidates, we can uniquely categorize events into boosted or resolved topologies. We demonstrate that SPA-Net accurately categorizes events in most cases for both $\mathrm{HH}$ and $\mathrm{HHH}$ events.
Overall, our results show that the boosted+resolved SPA-Net algorithm provides significant improvements in Higgs boson reconstruction across both $\mathrm{HHH}$ and $\mathrm{HH}$ events. Given the importance of maximizing the signal acceptance across different topologies to enhance the sensitivity to the small standard model cross sections, this method is a powerful tool for analyzing rare multi-Higgs boson final states at current and future colliders.
@@ Acknowledgements.
Appendix A Effect of $\mathrm{b}$-Tagging Efficiency on Reconstruction Performance
We further evaluated the impact of using an improved $\mathrm{b}$-tagging performance by emulating updated $\mathrm{b}$-tagging efficiencies based on recent experimental results CMS-DP-2024-066 . As shown in Fig. 15, the improved $\mathrm{b}$-tagging efficiency is about 85% instead of 70%, as used in previous sections.
Figure 15: The parametrized $\mathrm{b}$-tagging efficiencies of the state-of-the-art Unified Particle Transformer (UParT) and the Combined Secondary Vertex (CSV) algorithms. Markers indicate efficiencies calculated directly from the dataset, while the dotted lines represent the parameterized equations. The $\mathrm{b}$-tagging efficiency improves from approximately 70% to 85%.
This updated $\mathrm{b}$-tagging score was applied to both the $\chi^{2}$+BDT baseline method and as input to SPA-Net. We observe that the reconstruction efficiency and candidate purity improved for both methods under the updated $\mathrm{b}$-tagging configuration, as shown in Fig. 16.
Figure 16: Reconstruction purity (left) and efficiency (right) for the baseline and SPA-Net, shown for both the conservative $\mathrm{b}$ tagging (about 70% efficiency) and updated $\mathrm{b}$ tagging (about 85% efficiency). Both methods benefit similarly from the improved $\mathrm{b}$ tagging, indicating that the relative performance gain of SPA-Net remains stable across different $\mathrm{b}$-tagging configurations.
Notably, the relative improvements were comparable, indicating that the gains from enhanced $\mathrm{b}$ tagging are similar across both approaches and do not qualitatively affect the conclusions drawn from our comparison.